# Higgs-Confinement Transitions

### v2

@@ 3.4 Vafa-Witten Positivity
When the quark mass $m_{q}>0$ is positive, the QCD path integral measure can be regulated in a way that makes the path-integral measure positive definite [71, 72, 73].727272 Following these references, we work in the presence of a UV cutoff $\Lambda_{\text{UV}}$ and take $m_{q}$ to be the bare (rather than a renormalized or physical) quark mass. We refer to this as Vafa-Witten positivity. It has several important consequences:
@@
Note that ensuring Vafa-Witten positivity in the presence of background fields requires a particular choice of SPT counterterms.737373 This is in addition to the requirement that we set the theta-angle for the dynamical gauge fields to zero, $\theta_{c}=0$, as we have done throughout. A more invariant statement is that QCD is in the same SPT phase for any $m_{q}>0$. By choosing the counterterms in a particular way we can take it to be the trivial SPT, but this is not strictly necessary.
@@ 4.1 Defining the Theory
The presentation above (which applies for general $N$) obscures the full global symmetry of the $N=2$ theory we are considering here. Since bot $\psi^{a}$ and $\chi_{a}$ transform in the pseudo-real doublet representation of $SU(2)_{c}$, we can use the $SU(2)_{c}$ invariant symbols $\varepsilon^{ab},\varepsilon_{ab}$ to raise and lower the gauge indices,747474 We use the same conventions for $SU(2)$ index raising and lowering as for 2-component Weyl spinors. putting the left- and right-handed quarks on equal footing:
@@
A explained in section 1, theories with a gauged $(-1)^{F}$ symmetry (such as a conventional electronic superconductor) can never exhibit Higgs-confinement continuity, because Lorentz-scalar Higgs fields can at most break the gauge symmetry down to the ${\mathbb{Z}}_{2}^{F}$ subgroup generated by $(-1)^{F}$. Following [24], we therefore add a complex Lorentz-scalar Higgs field in the $SU(2)_{c}$ fundamental representation,
@@
Here we choose $i=1,2$ to be a doublet index under the $SU(2)_{f}$ flavor symmetry acting on the fermions.757575 A priori, the $SU(2)$ flavor symmetry acting on the Higgs field is decoupled from the $SU(2)_{f}$ symmetry acting on the fermions, but in section 4.2 below we add Yukawa couplings that identify them. Then $h^{a}_{i}$ transforms in the real, fundamental vector representation of the gauge and global $SO(4)_{c,f}$ group in (4.5). This is similar to the Higgs field in the standard model, with $SU(2)_{c}$ analogous to the weak gauge group and $SU(2)_{f}$ playing the role of the custodial symmetry. We also take $h^{a}_{i}$ to transform as a scalar under time-reversal,
@@ 4.2 SPT-Enforced Higgs-Confinement Transitions from Yukawa Couplings
The field content of our model does not allow for renormalizable Yukawa couplings, but it does have dimension-5 Yukawa couplings that preserve all gauge and global symmetries,767676 These couplings are reminiscent of the dimension-5 Weinberg operators in the standard model.
@@
where the extra factor of $i$ for the 4th component is due to Wick rotation. Here ${(\vec{\tau})^{a}}_{i}$ stands for the standard Pauli matrices ${(\tau^{I})^{a}}_{i}$, with $I=1,2,3$. We then make the change of variables777777 It can be checked that the matrix ${(\sigma_{E}^{I})^{a}}_{i}$ affecting the change of variables is an $SO(4)$ matrix, i.e. it is orthogonal of unit determinant.
@@
    Note that this is not the trivial SPT, because the flavor symmetry is $SO(3)_{f}^{\prime}=SU(2)_{2}^{\prime}/{\mathbb{Z}}_{2}$, so that the instanton number on spin manifolds can be half-integer,787878 Recall that the introduction of the Higgs field turned the previously bosonic $SU(2)_{c}$ QCD theory into a theory that requires a spin structure.
@@ 5.1 QCD with $N=3$ Colors and $N_{f}=3$ Flavors
Note that $PSU(3)$ symmetry allows more general background fields than $SU(3)$. In particular, the $PSU(3)$ instanton number can be fractional,797979 A $PSU(3)$ background gauge field with fractional instanton number can be thought of as an $SU(3)$ configuration with non-trivial ’t Hooft flux. even on spin manifolds, where
@@ 5.2 Higgs Fields, Yukawa Couplings, and Color-Flavor Locking
so that the entire Lagrangian is invariant under the gauge and flavor symmetries. Note that the Yukawa couplings preserve ${\mathsf{P}}$ and ${\mathsf{T}}$ (and hence also ${\mathsf{C}}$) as long as $y\in{\mathbb{R}}$. In fact we can say more: prior to adding the Yukawa coupling, the $U(1)_{B}$ symmetry acting on the fermions is not tied to the symmetry that rotates the Higgs field $h_{a}^{i}$ by an overall phase. We are thus free to use this symmetry to rotate $y$ by a phase, which is therefore of no physical consequence. Using this freedom, we take the Yukawa coupling constant to be positive,808080 The fact that the sign (and more generally the phase) of $y$ has no physical consequences distinguishes this model from the $SU(2)_{c}$ gauge theory example analyzed in section 4. There the presence of certain SPTs was directly tied to the signs of the Yukawas $y_{1},y_{2}$ in (4.13).
@@
Here the Higgs mass-squared $M_{h}^{2}\in{\mathbb{R}}$ is arbitrary real parameter, while $\lambda_{S},\lambda_{D}\in{\mathbb{R}}$ are single- and double-trace quartic couplings, respectively. The same class of scalar potentials were analyzed in [144, 145], where it was shown that imposing suitable bounds on $\lambda_{S},\lambda_{D}$ leads to a potential that is bounded from below and triggers a color-flavor locking (CFL) Higgs vev if we dial sufficiently deeply into the weakly-coupled Higgs phase by taking $M_{h}^{2}\ll-\Lambda_{\text{QCD}}^{2}$,818181 Specifically, we must impose (see appendix A of [144] and section 4.2 of [145]) $\lambda_{S}>0~{},\qquad\lambda_{S}+\lambda_{D}>0~{},\qquad\lambda_{S}+3\lambda_{D}>0~{}.$ (5.17) In principle $\lambda_{D}$ can have either sign; once the sign is fixed, either the second or the third inequality is redundant.
@@
By dialing the couplings $\lambda_{S,D}$, or higher-order terms in the potential, we could engineer other vevs for $h_{a}^{i}$.828282 For instance, we could take $\langle h_{a}^{i}\rangle$ to be of rank two, $\langle h_{a}^{i}\rangle=\text{diag}(v_{1},v_{2},0)~{},\qquad v_{1},v_{2}\neq 0~{},$ (5.19) which completely Higgses the $SU(3)_{c}$ gauge group without spontaneously breaking the $U(1)_{B}$ symmetry (since $\det(h_{a}^{i})=0$). It does however spontaneously break part of the $SU(3)_{f}$ flavor symmetry. The reason we focus on the CFL vev is that it characterizes the symmetry-breaking pattern that arises in high-density QCD, to be discussed in section 6.
@@ 5.3 SPTs and the Phase Diagram of $SU(3)$ Higgs-Yukawa-QCD
The determinant is $(4yv)^{2}-m_{q}^{2}$, with eigenvalues $4yv\pm m_{q}$. Thus the masses for singlet fermions are838383 Note that the eigenvectors are ${1\over\sqrt{2}}(\psi_{\bf 1}\pm\chi_{\bf 1})$. The change of basis matrix $U={1\over\sqrt{2}}\begin{pmatrix}1&1\\ -1&1\end{pmatrix}~{}.$ (5.27) is real orthogonal.
@@ 6 Three-Flavor QCD at Finite Baryon Density
  • Parity ${\mathsf{P}}$ and time-reversal ${\mathsf{T}}$ are preserved,848484 This is compatible with the broken ${\mathsf{C}}$ symmetry: since Lorentz symmetry is explicitly broken, the ${\mathsf{C}}{\mathsf{P}}{\mathsf{T}}$ theorem no longer holds. as are any flavor symmetries preserved by the quark masses, such as $U(1)_{B}$, and in the case of a common quark mass $m_{q}>0$, a vector-like $SU(3)_{f}$ (Gell-Mann) flavor symmetry. Spacetime translations and spatial rotations are also good symmetries.
@@
Another feature of turning on a real chemical potential $A_{0}=\mu_{B}$ in Lorentzian signature is that this turns into a purely imaginary background Wilson line $A^{E}_{4}=i\mu_{B}$ for the Wick rotated $U(1)_{B}$ background gauge field $A_{\mu}^{E}$ in Euclidean signature. Consequently the Euclidean path integral measure is no longer positive definite.858585 Although we only study the case of real chemical potential, it is also interesting to consider the case of purely imaginary chemical potential [146], which amounts to real $A_{4}^{E}$ and does not suffer from a sign problem (see also [147]). This has two important consequences:
@@ 6.1 Color-Flavor-Locking (CFL) at High Densities
Here we briefly review (following [94]) the behavior of three-flavor QCD at very large baryon chemical potential, $\mu_{B}\gg\Lambda_{\text{QCD}},m_{u},m_{d},m_{s}$. In this regime the differences between the quark masses can be largely ignored, and we will assume a common non-zero quark mass $m_{q}>0$, which preserves an $SU(3)_{f}$ flavor symmetry.868686 The story changes slightly if the quarks are exactly massless: in this case the high-density CFL phase has additional NGBs due to chiral symmetry breaking [150].
In the absence of interactions, the quarks form a degenerate Fermi gas with Fermi energy $\mu_{B}$. The light modes near the Fermi surface have large momenta $\sim\mu_{B}$, which implies that the effective gauge coupling $g(\mu_{B})\ll 1$ is very small thanks to asymptotic freedom. To study the BCS pairing instabilities of the Fermi surface, it suffices to consider one-gluon exchange between pairs of quarks. The most relevant attractive interaction is in the parity-even spin-0 (i.e. scalar rather than pseudo-scalar) channel, and it is anti-symmetric in both color and flavor indices.878787 Pairing in the spin-0 channel is favored because it is enhanced by the spherically symmetric Fermi surface. Since one-gluon exchange is attractive in the color anti-symmetric channel (and repulsive in the color-symmetric channel), restricting to spin-0 also implies anti-symmetry in flavor. Finally, the fact that the scalar channel is favored over the parity-odd pseudo-scalar one is not visible in perturbation theory and requires considerations involving QCD instantons. See section II.A of [94] for more details. This leads to a two-quark condensate with the following quantum numbers, or equivalently an expectation value for the following composite Higgs field [150],
@@
Since both models lead to weakly-coupled Higgs condensates with identical quantum numbers, many qualitative aspects of the physics are very similar:888888 Of course there are also differences, e.g. finite-density QCD is not Lorentz invariant and the sound speed in the high-density $U(1)_{B}$ superfluid phase is not the speed of light.
@@
    Note that the gap exponent scales as ${1/g}$, rather than the $1/g^{2}$ scaling familiar from standard BCS theory. This enhancement of the gap is due to the fact that the pairing interaction is long range, because it is mediated by gluons that are initially massless, before acquiring masses due to screening; by contrast, a short-range pairing interaction would lead to $1/g^{2}$. Comparing with section 5, where the fermion gap due to Higgsing is $\sim y|v|$, we see that the role of the dimensionless Yukawa coupling $y$ is played by $e^{-{K/g(\mu_{B})}}$.
@@ 6.2.1 Argument Based on Quark Pairing
    Higgs-Yukawa-QCD at $\mu_{B}=0$, but deep within its Higgs phase, pairs fermions in exactly the same channel (via the Yukawa couplings), leading to gaps that have exactly the same structure (including the relative factor of $2$ in (6.4)), though the overall magnitude of the gap is set by $y|v|$, rather then by BCS pairing. Once the fermions acquire these gaps, it can be checked explicitly that turning on a chemical potential $\mu_{B}$ deforms but never closes said gaps.898989 For a single free Dirac quark whose gap (equivalently, Majorana mass) $\Delta>0$ is much larger than its Dirac mass $m_{q}$ (so that $m_{q}$ can be ignored), the dispersion relation at 3-momentum $\vec{k}$ and chemical potential $\mu_{B}$ for the vector-like $U(1)_{B}$ symmetry acting on the quark is $\omega_{\pm}(\vec{k})=\sqrt{(|\vec{k}|\pm\mu_{B})^{2}+\Delta^{2}}~{},$ (6.6) see e.g. equation (30) in [94]. This shows that the fermions are never gapless – a conclusion that remains unchanged if we include $m_{q}$, as long as $\Delta\gtrsim m_{q}$.

### v3

@@ 3.4 Vafa-Witten Positivity
When the quark mass $m_{q}>0$ is positive, the QCD path integral measure can be regulated in a way that makes the path-integral measure positive definite [71, 72, 73].747474 Following these references, we work in the presence of a UV cutoff $\Lambda_{\text{UV}}$ and take $m_{q}$ to be the bare (rather than a renormalized or physical) quark mass. We refer to this as Vafa-Witten positivity. It has several important consequences:
@@
Note that ensuring Vafa-Witten positivity in the presence of background fields requires a particular choice of SPT counterterms.757575 This is in addition to the requirement that we set the theta-angle for the dynamical gauge fields to zero, $\theta_{c}=0$, as we have done throughout. A more invariant statement is that QCD is in the same SPT phase for any $m_{q}>0$. By choosing the counterterms in a particular way we can take it to be the trivial SPT, but this is not strictly necessary.
@@ 4.1 Defining the Theory
The presentation above (which applies for general $N$) obscures the full global symmetry of the $N=2$ theory we are considering here. Since bot $\psi^{a}$ and $\chi_{a}$ transform in the pseudo-real doublet representation of $SU(2)_{c}$, we can use the $SU(2)_{c}$ invariant symbols $\varepsilon^{ab},\varepsilon_{ab}$ to raise and lower the gauge indices,767676 We use the same conventions for $SU(2)$ index raising and lowering as for 2-component Weyl spinors. putting the left- and right-handed quarks on equal footing:
@@
A explained in section 1, theories with a gauged $(-1)^{F}$ symmetry (such as a conventional electronic superconductor) can never exhibit Higgs-confinement continuity, because Lorentz-scalar Higgs fields can at most break the gauge symmetry down to the ${\mathbb{Z}}_{2}^{F}$ subgroup generated by $(-1)^{F}$. Following [24], we therefore add a complex Lorentz-scalar Higgs field in the $SU(2)_{c}$ fundamental representation, so that $(-1)^{F}$ is no longer gauged and we can completely break the gauge group,
@@
Here we choose $i=1,2$ to be a doublet index under the $SU(2)_{f}$ flavor symmetry acting on the fermions.777777 A priori, the $SU(2)$ flavor symmetry acting on the Higgs field is decoupled from the $SU(2)_{f}$ symmetry acting on the fermions, but in section 4.2 below we add Yukawa couplings that identify them. Then $h^{a}_{i}$ transforms in the real, fundamental vector representation of the gauge and global $SO(4)_{c,f}$ group in (4.5). This is similar to the Higgs field in the standard model, with $SU(2)_{c}$ analogous to the weak gauge group and $SU(2)_{f}$ playing the role of the custodial symmetry. We also take $h^{a}_{i}$ to transform as a scalar under time-reversal,
@@ 4.2 SPT-Enforced Higgs-Confinement Transitions from Yukawa Couplings
The field content of our model does not allow for renormalizable Yukawa couplings, but it does have dimension-5 Yukawa couplings that preserve all gauge and global symmetries,787878 These couplings are reminiscent of the dimension-5 Weinberg operators in the standard model.
@@
where the extra factor of $i$ for the 4th component is due to Wick rotation. Here ${(\vec{\tau})^{a}}_{i}$ stands for the standard Pauli matrices ${(\tau^{I})^{a}}_{i}$, with $I=1,2,3$. We then make the change of variables797979 It can be checked that the matrix ${(\sigma_{E}^{I})^{a}}_{i}$ affecting the change of variables is an $SO(4)$ matrix, i.e. it is orthogonal of unit determinant.
@@
    Note that this is not the trivial SPT, because the flavor symmetry is $SO(3)_{f}^{\prime}=SU(2)_{2}^{\prime}/{\mathbb{Z}}_{2}$, so that the instanton number on spin manifolds can be half-integer,808080 Recall that the introduction of the Higgs field turned the previously bosonic $SU(2)_{c}$ QCD theory into a theory that requires a spin structure.
@@ 5.1 QCD with $N=3$ Colors and $N_{f}=3$ Flavors
Note that $PSU(3)$ symmetry allows more general background fields than $SU(3)$. In particular, the $PSU(3)$ instanton number can be fractional,818181 A $PSU(3)$ background gauge field with fractional instanton number can be thought of as an $SU(3)$ configuration with non-trivial ’t Hooft flux. even on spin manifolds, where
@@ 5.2 Higgs Fields, Yukawa Couplings, and Color-Flavor Locking
so that the entire Lagrangian is invariant under the gauge and flavor symmetries. Note that the Yukawa couplings preserve ${\mathsf{P}}$ and ${\mathsf{T}}$ (and hence also ${\mathsf{C}}$) as long as $y\in{\mathbb{R}}$. In fact we can say more: prior to adding the Yukawa coupling, the $U(1)_{B}$ symmetry acting on the fermions is not tied to the symmetry that rotates the Higgs field $h_{a}^{i}$ by an overall phase. We are thus free to use this symmetry to rotate $y$ by a phase, which is therefore of no physical consequence. Using this freedom, we take the Yukawa coupling constant to be positive,828282 The fact that the sign (and more generally the phase) of $y$ has no physical consequences distinguishes this model from the $SU(2)_{c}$ gauge theory example analyzed in section 4. There the presence of certain SPTs was directly tied to the signs of the Yukawas $y_{1},y_{2}$ in (4.13).
@@
Here the Higgs mass-squared $M_{h}^{2}\in{\mathbb{R}}$ is arbitrary real parameter, while $\lambda_{S},\lambda_{D}\in{\mathbb{R}}$ are single- and double-trace quartic couplings, respectively. The same class of scalar potentials were analyzed in [145, 146], where it was shown that imposing suitable bounds on $\lambda_{S},\lambda_{D}$ leads to a potential that is bounded from below and triggers a color-flavor locking (CFL) Higgs vev if we dial sufficiently deeply into the weakly-coupled Higgs phase by taking $M_{h}^{2}\ll-\Lambda_{\text{QCD}}^{2}$,838383 Specifically, we must impose (see appendix A of [145] and section 4.2 of [146]) $\lambda_{S}>0~{},\qquad\lambda_{S}+\lambda_{D}>0~{},\qquad\lambda_{S}+3\lambda_{D}>0~{}.$ (5.17) In principle $\lambda_{D}$ can have either sign; once the sign is fixed, either the second or the third inequality is redundant.
@@
By dialing the couplings $\lambda_{S,D}$, or higher-order terms in the potential, we could engineer other vevs for $h_{a}^{i}$.848484 For instance, we could take $\langle h_{a}^{i}\rangle$ to be of rank two, $\langle h_{a}^{i}\rangle=\text{diag}(v_{1},v_{2},0)~{},\qquad v_{1},v_{2}\neq 0~{},$ (5.19) which completely Higgses the $SU(3)_{c}$ gauge group without spontaneously breaking the $U(1)_{B}$ symmetry (since $\det(h_{a}^{i})=0$). It does however spontaneously break part of the $SU(3)_{f}$ flavor symmetry. The reason we focus on the CFL vev is that it characterizes the symmetry-breaking pattern that arises in high-density QCD, to be discussed in section 6.
@@ 5.3 SPTs and the Phase Diagram of $SU(3)$ Higgs-Yukawa-QCD
The determinant is $(4yv)^{2}-m_{q}^{2}$, with eigenvalues $4yv\pm m_{q}$. Thus the masses for singlet fermions are858585 Note that the eigenvectors are ${1\over\sqrt{2}}(\psi_{\bf 1}\pm\chi_{\bf 1})$. The change of basis matrix $U={1\over\sqrt{2}}\begin{pmatrix}1&1\\ -1&1\end{pmatrix}~{}.$ (5.27) is real orthogonal.
@@ 6 Three-Flavor QCD at Finite Baryon Density
  • Parity ${\mathsf{P}}$ and time-reversal ${\mathsf{T}}$ are preserved,868686 This is compatible with the broken ${\mathsf{C}}$ symmetry: since Lorentz symmetry is explicitly broken, the ${\mathsf{C}}{\mathsf{P}}{\mathsf{T}}$ theorem no longer holds. as are any flavor symmetries preserved by the quark masses, such as $U(1)_{B}$, and in the case of a common quark mass $m_{q}>0$, a vector-like $SU(3)_{f}$ (Gell-Mann) flavor symmetry. Spacetime translations and spatial rotations are also good symmetries.
@@
Another feature of turning on a real chemical potential $A_{0}=\mu_{B}$ in Lorentzian signature is that this turns into a purely imaginary background Wilson line $A^{E}_{4}=i\mu_{B}$ for the Wick rotated $U(1)_{B}$ background gauge field $A_{\mu}^{E}$ in Euclidean signature. Consequently the Euclidean path integral measure is no longer positive definite.878787 Although we only study the case of real chemical potential, it is also interesting to consider the case of purely imaginary chemical potential [147], which amounts to real $A_{4}^{E}$ and does not suffer from a sign problem (see also [148]). This has two important consequences:
@@ 6.1 Color-Flavor-Locking (CFL) at High Densities
Here we briefly review (following [95]) the behavior of three-flavor QCD at very large baryon chemical potential, $\mu_{B}\gg\Lambda_{\text{QCD}},m_{u},m_{d},m_{s}$. In this regime the differences between the quark masses can be largely ignored, and we will assume a common non-zero quark mass $m_{q}>0$, which preserves an $SU(3)_{f}$ flavor symmetry.888888 The story changes slightly if the quarks are exactly massless: in this case the high-density CFL phase has additional NGBs due to chiral symmetry breaking [151].
In the absence of interactions, the quarks form a degenerate Fermi gas with Fermi energy $\mu_{B}$. The light modes near the Fermi surface have large momenta $\sim\mu_{B}$, which implies that the effective gauge coupling $g(\mu_{B})\ll 1$ is very small thanks to asymptotic freedom. To study the BCS pairing instabilities of the Fermi surface, it suffices to consider one-gluon exchange between pairs of quarks. The most relevant attractive interaction is in the parity-even spin-0 (i.e. scalar rather than pseudo-scalar) channel, and it is anti-symmetric in both color and flavor indices.898989 Pairing in the spin-0 channel is favored because it is enhanced by the spherically symmetric Fermi surface. Since one-gluon exchange is attractive in the color anti-symmetric channel (and repulsive in the color-symmetric channel), restricting to spin-0 also implies anti-symmetry in flavor. Finally, the fact that the scalar channel is favored over the parity-odd pseudo-scalar one is not visible in perturbation theory and requires considerations involving QCD instantons. See section II.A of [95] for more details. This leads to a two-quark condensate with the following quantum numbers, or equivalently an expectation value for the following composite Higgs field [151],
@@
Since both models lead to weakly-coupled Higgs condensates with identical quantum numbers, many qualitative aspects of the physics are very similar:909090 Of course there are also differences, e.g. finite-density QCD is not Lorentz invariant and the sound speed in the high-density $U(1)_{B}$ superfluid phase is not the speed of light.
@@
    Note that the gap exponent scales as ${1/g}$, rather than the $1/g^{2}$ scaling familiar from standard BCS theory. This enhancement of the gap is due to the fact that the pairing interaction is long range, because it is mediated by gluons that are initially massless, before acquiring masses due to screening; by contrast, a short-range pairing interaction would lead to $1/g^{2}$. Comparing with section 5, where the fermion gap due to Higgsing is $\sim y|v|$, we see that the role of the dimensionless Yukawa coupling $y$ is played by $e^{-{K/g(\mu_{B})}}$.919191 As mentioned at the beginning of section 6.1, we focus on the regime where the chemical potential $\mu_{B}$ is much larger than the quark masses – sufficiently large so that $\Delta_{\bf 1},\Delta_{\bf 8}\gg m_{u},m_{d},m_{s}$.
@@ 6.2.1 Argument Based on Quark Pairing
    Higgs-Yukawa-QCD at $\mu_{B}=0$, but deep within its Higgs phase, pairs fermions in exactly the same channel (via the Yukawa couplings), leading to gaps that have exactly the same structure (including the relative factor of $2$ in (6.4)), though the overall magnitude of the gap is set by $y|v|$, rather then by BCS pairing. Once the fermions acquire these gaps, it can be checked explicitly that turning on a chemical potential $\mu_{B}$ deforms but never closes said gaps.929292 For a single free Dirac quark whose gap (equivalently, Majorana mass) $\Delta>0$ is much larger than its Dirac mass $m_{q}$ (so that $m_{q}$ can be ignored), the dispersion relation at 3-momentum $\vec{k}$ and chemical potential $\mu_{B}$ for the vector-like $U(1)_{B}$ symmetry acting on the quark is $\omega_{\pm}(\vec{k})=\sqrt{(|\vec{k}|\pm\mu_{B})^{2}+\Delta^{2}}~{},$ (6.6) see e.g. equation (30) in [95]. This shows that the fermions are never gapless – a conclusion that remains unchanged if we include $m_{q}$, as long as $\Delta\gtrsim m_{q}$.